# Writing - main part of original draft, Visualization, Software, Methodology, Derivation of error bounds, Verification, Conceptualization

### v1

@@ Abstract
Fiber orientation is decisive for the mechanical properties and thus for the performance of composite materials. During manufacturing, variations in material and process parameters can significantly influence the exact fiber orientation. We employ multilevel polynomial surrogates to model the propagation of uncertain material properties in the injection molding process. To ensure reliable uncertainty quantification, a key focus is deriving novel error bounds for statistical measures of a quantity of interest, computed via these surrogates. To verify these bounds, we conduct numerical experiments using the Cross-WLF viscosity model alongside the Hagen-Poiseuille flow in a rectangular channel. In particular, the impact of uncertainties in fiber length and matrix temperature on the fractional anisotropy of fiber orientation is investigated. The Folgar-Tucker equation and the improved anisotropic rotary diffusion model are used, incorporating recently established analytical solutions of these models as part of our verification. Our results demonstrate that the investigated method significantly improves upon standard Monte Carlo estimation, while also providing error guarantees. These findings offer the first step toward a reliable and practical tool for optimizing fiber-reinforced polymer manufacturing processes in the future.
@@ 2.3 Fiber orientation modeling
$\begin{split}\mathcal{S}_{+}=\bigl{\{}\mathbf{A}\in\mathbb{R}^{3\times 3}:\,\,&\mathbf{A}=\mathbf{A}^{\top}\\ &x^{\top}\mathbf{A}x\geq 0\,\,\,\forall x\in\mathbb{R}^{3},\\ &\text{tr}(\mathbf{A})=1\bigr{\}}\\ \end{split}$ (4)
@@
$\begin{split}\dot{\mathbf{A}}&=\mathbf{W}\mathbf{A}-\mathbf{A}\mathbf{W}+\xi\left(\mathbf{D}\mathbf{A}+\mathbf{A}\mathbf{D}-2\mathbf{D}:\mathbb{A}\right)\\ &+2C_{I}\dot{\gamma}(\mathbf{I}-3\mathbf{A})\end{split}$ (6)
@@
$\begin{split}\dot{\mathbf{A}}&=\mathbf{W}\mathbf{A}-\mathbf{A}\mathbf{W}+\xi\left(\mathbf{D}\mathbf{A}+\mathbf{A}\mathbf{D}-2\mathbf{D}:\mathbb{A}\right)\\ &+\dot{\gamma}\left[2\mathbf{C}-2\text{tr}(\mathbf{C})\mathbf{A}-5(\mathbf{C}\mathbf{A}+\mathbf{A}\mathbf{C})+10\mathbb{A}:\mathbf{C}\right].\end{split}$ (7)
@@ 3.3 Error bounds for statistical quantities
where $f_{Q}$ is the PDF of $Q$ and $\|f\|_{\infty}:=\sup_{x\in\mathbb{R}}|f(x)|$ denotes the supremum norm. Finally, for the $p$-quantile with probability $p\in(0,1)$ of a random variable $X$, that is
$q_{X}:=\inf\{q\in\mathbb{R}:F_{X}(q)\geq p\},$ (31)
we have
@@ 4 Numerical experiments
Figure 2: Hagen-Poiseuille flow $v(x,y)$, Eq. (1), with viscosity $\eta=\eta(T)$ for $T=565$ K. The red dot marks the arbitrary selected point in the channel $(x_{0},y_{0})$ and the arrows show the gradient.
@@ 4.1 Fixed material point
Figure 3: Response surface $Q=F\left(\mathbf{A}[t_{\text{end}};\cdot]\right)$ for the parameter domains $\Omega_{T}=[550,580]$ K and $\Omega_{L_{f}}=[0.38\pm 50\%]$ mm, computed using the FTE and iARD model as FOM, respectively, with isotropic initial orientation tensor $\mathbf{A}(0)=\mathbf{I}/3$.
@@ 4.1.2 Reference values for exact statistics
To compute the errors needed for verification of the bounds from Section 3.3, we utilize the analytical solutions from Appendix A. The eigenvalues required for evaluating the fractional anisotropy, Eq. (41), are computed numerically. Our experiments demonstrate that the error introduced by these computations is negligible compared to the error inherent in solving the FOM itself. As the QoI statistics are not analytically available, they must be estimated. For this purpose, we utilize a quasi-Monte Carlo estimation on a Sobol sequence
@@ 4.2 Results and discussion for fixed material point
Figure 4: Input tolerance $\epsilon$ vs. the resulting error $\|Q-\hat{Q}_{\epsilon}\|$. Also shown are the errors for the expected value $\mathbb{E}[Q]$, standard deviation $\mathbb{V}[Q]^{1/2}$, CDF $F_{Q}$, and the $p$-quantile $q_{Q}$. The left plot presents results based on the FTE, while the right plot corresponds to the iARD model.
@@
is approximated by the maximum difference evaluated at the points of the Sobol sample $\omega_{S}$. For the FTE, the result is depicted in the left plot of Figure 4, which shows that
@@
Table 1: Comparison of expected value $\mathbb{E}[Q]$, standard deviation $\mathbb{V}[Q]^{1/2}$, CDF $F_{Q}$ at the arbitrarily chosen point $0.953$, and the $p$-quantile $q_{Q}$ with $p=0.99$ for the FTE and iARD model, including their reference values and relative errors. The relative errors were computed using the polynomial surrogate $\hat{Q}_{\epsilon}$ with $\epsilon=10^{-3}$.
This can also be observed in Figure 5, which depicts the probability density function (PDF) and histogram for both models, computed using a kernel density estimation (KDE) [36] of their surrogates on the Sobol sample $\omega_{S}$. The iARD model’s density function exhibits a broader support, indicating a higher variance. Similar to what can be observed in Figure 3, the iARD model predicts a higher fractional anisotropy.
Figure 5: Shown are the PDFs and histograms of the QoI, fractional anisotropy, for the FTE and the iARD model. They were computed using a KDE of their polynomial surrogates $\hat{Q}_{\epsilon}$ with $\epsilon=10^{-3}$ on the Sobol sample $\omega_{S}$.
@@ 4.3 Entire flow channel
(c) Standard deviation $\mathbb{V}[\hat{Q}_{(x,y)}]^{1/2}$
(d) Probability $\mathbb{P}(0.85\leq\hat{Q}_{(x,y)}\leq 0.9)$
Figure 6: Hagen-Poiseuille flow (6(a)) and surrogate statistics of the fractional anisotropy, $\hat{Q}_{(x,y)}=F(A)$, according to Eq. (41), across the channel (6(b) - 6(d)). All computations were performed using the FTE, Eq. (6).
The expectations and variances are computed exactly by post-processing the Legendre coefficients, as described in Section 3.2.
4.4 Results and discussion for the entire flow channel
It can be observed that the expected fractional anisotropy spans almost the entire range from 00 to $1$ across the channel. This is primarily influenced by the Hagen-Poiseuille flow and its associated velocity gradient, Eq. (42). For example, the horizontal middle strip of the channel exhibits low anisotropy due to the small values in the velocity gradient, which do not significantly reorient the fibers (see Figure 6(a)). This effect is also noticeable in the corners of the channel. Figure 6(c) depicts the standard deviation within the channel. The QoI, fractional anisotropy, is not significantly impacted by uncertain temperature or fiber length in areas with sufficiently high velocity gradient, resulting in low standard deviation. Conversely, in regions where the gradient is almost zero, fiber reorientation is minimal, so uncertainties in these parameters also have little effect on the anisotropy, as observed in the slim horizontal middle strip of the channel. Of particular interest is the ring surrounding this middle strip, which exhibits the highest standard deviation. In this area, the material properties, temperature and fiber length, have the most pronounced influence on the QoIs variability. This region likely represents a zone where the underlying velocity gradient is neither too low nor too high, allowing the effects of uncertainties in material properties to become more apparent. Finally, Figure 6(d) depicts the CDF evaluation at arbitrarily chosen points. Specifically, we examine the probability that the fractional anisotropy lies between $0.85$ and $0.9$, which can be computed by
$\mathbb{P}(0.85\leq\hat{Q}_{(x,y)}\leq 0.9)=F_{\hat{Q}_{(x,y)}}(0.9)-F_{\hat{Q}_{(x,y)}}(0.85),$ (54)
where $F_{\hat{Q}_{(x,y)}}$ is the CDF of the surrogates partial evaluation. Gradual oscillations can be observed, which likely result from the polynomial approximation of the response surface. Similar to the standard deviation, a ring or contour line is present, within which the considered values of the QoI can be found, supported by probabilistic information.
5 Conclusion
This study focuses on modeling fiber orientation in composite molding processes using novel error bounds for statistics of a quantity of interest (QoI) computed via polynomial surrogates. These bounds are based on the $L^{2}$ error of the polynomial approximation. Numerical experiments employing the Cross-WLF viscosity model, Hagen-Poiseuille flow, the Folgar-Tucker equation (FTE), and the improved anisotropic rotary diffusion (iARD) model verify the proposed theory. Our findings show that uncertainties in material properties, such as temperature and fiber length, have significant impact on the fractional anisotropy of fiber orientation. We observe that both the FTE and the iARD model produce similar results at first glance for deterministic simulations. However, with polynomial surrogates, it is possible to reveal subtle differences between these fiber orientation models, which highlights another aspect of their usefulness.
In addition to investigating the fiber orientation distribution in a single material point, we extended the approach to learn the spatial distribution of the fractional anisotropy across the entire rectangular channel domain. To support this, theoretical bounds for the $L^{2}$ error control of statistics over the entire domain through partial evaluations are presented. These considerations indicate that while the derivation of a global, geometry-wide mapping of the QoI to pointwise evaluations is limited from a theoretical point of view, it is still effective for qualitative investigations of fiber orientation. One limitation of the presented approach is its reliance on finite-dimensional uncertainties, instead of, for example, time-varying random fields. Additionally, the numerical experiments are conducted on relatively simple geometries with smooth response surfaces, where polynomial surrogates perform well. Their performance may drop in more complicated settings with non-smooth responses.
Despite these limitations, the derived error bounds are not inherently tied to polynomial surrogates and could extend to alternative methods, such as neural networks. Nonetheless, the multilevel polynomial least squares approach retains a distinct advantage through its a-priori $L^{2}$ error bounds, providing a strong theoretical foundation. Overall, this research contributes to the understanding and application of polynomial surrogates and provides a robust framework for uncertainty quantification, error control and model comparison in the context of fiber orientation.
Appendix A Analytical solution to fiber orientation models
We utilize the analytical solution derived by Winters et al. [27] and present the exact formulas needed to perform the numerical experiments in this work. Let $\mathbf{e}=\mathbf{e}_{1}+\mathbf{e}_{5}+\mathbf{e}_{9}$, where $\mathbf{e}_{i}\in\mathbb{R}^{9}$ denotes the unit vector with a 1 in the $i$-th position and 0 elsewhere. Further, let $\texttt{vec}:\mathbb{R}^{3\times 3}\rightarrow\mathbb{R}^{9}$ be the mapping for the column-wise vectorization of a matrix into a vector, and $\texttt{mat}:\mathbb{R}^{9}\rightarrow\mathbb{R}^{3\times 3}$ the reverse operation. The following analytical solutions are valid for a hybrid closure [38], defined as
$\mathbb{A}^{H}=(1-f)\mathbb{A}^{L}+f\mathbb{A}^{Q},$ (55)
where $f\in[0,1]$, and $\mathbb{A}^{Q}_{ijkl}=\mathbf{A}_{ij}\mathbf{A}_{kl}$ represents the quadratic closure [34], and
$\begin{split}\mathbb{A}^{L}_{ijkl}&=-\frac{1}{35}\left(\delta_{ij}\delta_{kl}\!+\!\delta_{ik}\delta_{jl}\!+\!\delta_{il}\delta_{jk}\right)\\ &+\frac{1}{7}\left(\mathbf{A}_{ij}\delta_{kl}\!+\!\mathbf{A}_{ik}\delta_{jl}\!+\!\mathbf{A}_{il}\delta_{jk}\!+\!\mathbf{A}_{jl}\delta_{ik}\!+\!\mathbf{A}_{jk}\delta_{il}\!+\!\mathbf{A}_{kl}\delta_{ij}\right)\end{split}$ (56)
denotes the linear closure [39]. Note that $f$ must not depend on $\mathbf{A}$, which excludes the classical hybrid closure with $f=1-27\,\text{det}(\mathbf{A})$. Define the variables $\mathbf{L_{\xi}}:=\mathbf{W}+\xi\mathbf{D}$ and $\hat{f}:=2(f-1)$, as well as $\mathcal{L}:\mathbb{R}^{3\times 3}\rightarrow\mathbb{R}^{9\times 9}$ with
$\begin{split}\mathcal{L}(\mathbf{A})&=\frac{1}{7}\text{tr}(\mathbf{A})\,\mathbf{I}+\frac{2}{7}(\mathbf{I}\otimes\mathbf{A}+\mathbf{A}\otimes\mathbf{I})\\ &+\frac{1}{7}\mathbf{e}\,\texttt{vec}(\mathbf{A})^{\top}-\frac{2}{35}\texttt{vec}(\mathbf{A})\,\mathbf{e}^{\top}-\frac{1}{35}\text{tr}(\mathbf{A})\,\mathbf{e}\mathbf{e}^{\top}.\end{split}$ (57)
Using the notation introduced above, the solution to the Folgar-Tucker equation, Eq. (6), with initial orientation tensor $\mathbf{A}_{0}\in\mathbb{R}^{3\times 3}$ is given by
$\mathbf{A}(t)=\texttt{mat}\left(e^{t\mathbf{M}}\,\texttt{vec}\left(\mathbf{A}_{0}\right)\right)$ (58)
with
$\mathbf{M}=\mathbf{I}\otimes\mathbf{L_{\xi}}+\mathbf{L_{\xi}}\otimes\mathbf{I}+2C_{I}\dot{\gamma}\left(\mathbf{e}\mathbf{e}^{\top}-3\mathbf{I}\right)+\xi\hat{f}\mathcal{L}(\mathbf{D}).$ (59)
Similarly, defining $\mathbf{E}:=\xi\mathbf{D}-5\dot{\gamma}\mathbf{C}$ as suggested by Favaloro and Tucker [40], the solution to the anisotropic rotary diffusion model, Eq. (7), with initial orientation tensor $\mathbf{A}_{0}\in\mathbb{R}^{3\times 3}$ is given by
$\mathbf{A}(t)=\texttt{mat}\left(e^{t\mathbf{M}}\,\texttt{vec}\left(\mathbf{A}_{0}\right)\right)$ (60)
with
$\begin{split}\mathbf{M}&=\mathbf{I}\otimes\mathbf{L_{\xi}}+\mathbf{L_{\xi}}\otimes\mathbf{I}+2\dot{\gamma}\left(\texttt{vec}(\mathbf{C})\,\mathbf{e}^{\top}-\text{tr}(\mathbf{C})\,\mathbf{I}\right)\\ &-5\dot{\gamma}(\mathbf{I}\otimes\mathbf{C}+\mathbf{C}\otimes\mathbf{I})+\hat{f}\mathcal{L}(\mathbf{E}).\end{split}$ (61)
Appendix B Error bounds for statistical quantities
In the following sections, we provide a detailed explanation on the derivation of error bounds for several statistics that are approximated using a surrogate $\hat{Q}$. The aim in each case is to ensure that the upper bound includes the $L^{2}$ approximation error $\|Q-\hat{Q}\|$ of the surrogate.
B.1 Standard deviation
Using the reverse triangle inequality it follows for the difference of the standard deviations
$\begin{split}\left|\mathbb{V}[Q]^{1/2}-\mathbb{V}[\hat{Q}]^{1/2}\right|^{2}&=\left|\|Q-\mathbb{E}[Q]\|-\|\hat{Q}-\mathbb{E}[\hat{Q}]\|\right|^{2}\\ &\leq\|(Q-\hat{Q})-(\mathbb{E}[Q]-\mathbb{E}[\hat{Q}])\|^{2}\\ &=\|Q-\hat{Q}\|^{2}+\|\mathbb{E}[Q]-\mathbb{E}[\hat{Q}]\|^{2}\\ &\quad-2\langle Q-\hat{Q},\,\mathbb{E}[Q]-\mathbb{E}[\hat{Q}]\rangle\\ &=\|Q-\hat{Q}\|^{2}-|\mathbb{E}[Q]-\mathbb{E}[\hat{Q}]|^{2}\\ &\leq\|Q-\hat{Q}\|^{2},\end{split}$ (62)
where $\langle f,g\rangle=\int_{\Omega}f(\omega)g(\omega)\,\mu(d\omega)$ denotes the $L^{2}_{\mu}$ inner product for elements $f,g\in L^{2}_{\mu}(\Omega)$.
B.2 Cumulative distribution function
In addition to the $L^{2}$ norm considered so far, we define, for $p\in(0,\infty)$, the $L^{p}$ norm as
$\|f\|_{L^{p}_{\mu}(\Omega)}=\left(\int_{\Omega}|f(\omega)|^{p}\,\mu(d\omega)\right)^{\frac{1}{p}}$ (63)
and abbreviate $\|\cdot\|_{L^{p}}:=\|\cdot\|_{L^{p}_{\mu}(\Omega)}$.
Lemma 1
Let $X$ and $Y$ be random variables on the same probability space. Then, if $X$ has a bounded density $f_{X}$, it holds for every $0<p<\infty$ that
$\begin{split}\sup_{x\in\mathbb{R}}\mathbb{E}\bigl{[}|\mathbbm{1}_{\{X\leq x\}}-\mathbbm{1}&{}_{\{Y\leq x\}}|\bigr{]}\\ &\leq 3\left(\|f_{X}\|_{\infty}\mathbb{E}\bigl{[}|X-Y|^{p}\bigr{]}^{\frac{1}{p}}\right)^{\frac{p}{p+1}}\end{split}$ (64)

### v2

# Methodology, Conceptualization, Resources, Writing - review and editing

@@ Abstract
Fiber orientation is decisive for the mechanical performance of composite materials. During manufacturing, variations in material and process parameters can influence fiber orientation. We employ multilevel polynomial surrogates to model the propagation of uncertain material properties in the injection molding process. To ensure reliable uncertainty quantification, a key focus is deriving novel error bounds for statistical measures of a quantity of interest. Numerical experiments employ the Cross-WLF viscosity model and Hagen-Poiseuille flow to investigate the impact of uncertainties in fiber length and matrix temperature on the fractional anisotropy of fiber orientation. The Folgar-Tucker equation and the improved anisotropic rotary diffusion model, incorporating analytical solutions, are used for verification. Results show that the method improves significantly upon standard Monte Carlo estimation, while also providing error guarantees. These findings offer the first step toward a reliable and practical tool for optimizing fiber-reinforced polymer manufacturing processes in the future.
@@ 2.3 Fiber orientation modeling
$\mathcal{S}_{+}=\bigl{\{}\mathbf{A}\in\mathbb{R}^{3\times 3}:\mathbf{A}=\mathbf{A}^{\top},\quad x^{\top}\mathbf{A}x\geq 0\,\,\,\forall x\in\mathbb{R}^{3},\quad\text{tr}(\mathbf{A})=1\bigr{\}}$ (4)
@@
$\dot{\mathbf{A}}=\mathbf{W}\mathbf{A}-\mathbf{A}\mathbf{W}+\xi\left(\mathbf{D}\mathbf{A}+\mathbf{A}\mathbf{D}-2\mathbf{D}:\mathbb{A}\right)+2C_{I}\dot{\gamma}(\mathbf{I}-3\mathbf{A})$ (6)
@@
$\dot{\mathbf{A}}=\mathbf{W}\mathbf{A}-\mathbf{A}\mathbf{W}+\xi\left(\mathbf{D}\mathbf{A}+\mathbf{A}\mathbf{D}-2\mathbf{D}:\mathbb{A}\right)+\dot{\gamma}\left[2\mathbf{C}-2\text{tr}(\mathbf{C})\mathbf{A}-5(\mathbf{C}\mathbf{A}+\mathbf{A}\mathbf{C})+10\mathbb{A}:\mathbf{C}\right].$ (7)
@@ 3.3 Error bounds for statistical quantities
where $f_{Q}$ is the PDF of $Q$ and $\|f\|_{\infty}:=\sup_{x\in\mathbb{R}}|f(x)|$ denotes the supremum norm. Finally, for the $p$-quantile with probability $p\in(0,1)$ of a random variable $X$, that is $q_{X}:=\inf\{q\in\mathbb{R}:F_{X}(q)\geq p\}$, we have
@@ 4 Numerical experiments
[…]
@@ 4.1 Fixed material point
Figure 3: Response surface $Q=F\left(\mathbf{A}[t_{\text{end}};\cdot]\right)$ showing the fractional anisotropy over the parameter domains $\Omega_{T}=[550,580]$ K (temperature) and $\Omega_{L_{f}}=[0.38\pm 50\%]$ mm (fiber length). Computations use the FTE and iARD model as FOM, respectively, with isotropic initial orientation tensor $\mathbf{A}(0)=\mathbf{I}/3$.
@@ 4.1.2 Reference values for exact statistics
To compute the errors needed for verification of the bounds from Section 3.3, we utilize the analytical solutions from Appendix A. The eigenvalues required for evaluating the fractional anisotropy, Eq. (40), are computed numerically. Our experiments demonstrate that the error introduced by these computations is negligible compared to the error inherent in solving the FOM itself. As the QoI statistics are not analytically available, they must be estimated. For this purpose, we utilize a quasi-Monte Carlo estimation on the Sobol sequence
@@ 4.2 Results and discussion for fixed material point
Figure 4: Input tolerance $\epsilon$ vs. the resulting error $\|Q-\hat{Q}_{\epsilon}\|$ of the polynomial approximation of fractional anisotropy $Q=F(\mathbf{A})$. Also shown are the errors for the expected value $\mathbb{E}[Q]$, standard deviation $\mathbb{V}[Q]^{1/2}$, CDF $F_{Q}$, and the $p$-quantile $q_{Q}$. The left plot presents results based on the FTE, while the right plot corresponds to the iARD model.
@@
is approximated by the maximum difference evaluated at the points of the Sobol sample $\omega_{S}$, Eq. (48). For the FTE, the result is depicted in the left plot of Figure 4, which shows that
@@
Table 1: Comparison of expected value $\mathbb{E}[Q]$, standard deviation $\mathbb{V}[Q]^{1/2}$, CDF $F_{Q}$ at the arbitrarily chosen point $0.953$, and the $p$-quantile $q_{Q}$ with $p=0.99$ for the fractional anisotropy $Q=F(\mathbf{A})$ using the FTE and iARD model, including their reference values and relative errors. The relative errors were computed using the polynomial surrogate $\hat{Q}_{\epsilon}$ with $\epsilon=10^{-3}$.
This can also be observed in Figure 5, which depicts the probability density function (PDF) and histogram for both models, computed using a kernel density estimation (KDE) [36] of the evaluation of their surrogates on the Sobol sample $\omega_{S}$, Eq. (48). The iARD model’s density function exhibits a broader support, indicating a higher variance. Similar to what can be observed in Figure 3, the iARD model predicts a higher fractional anisotropy.
Figure 5: Shown are the PDFs and histograms of the QoI, fractional anisotropy, for the FTE and the iARD model. They were computed using a KDE of evaluations of their polynomial surrogates $\hat{Q}_{\epsilon}$ with $\epsilon=10^{-3}$ on the Sobol sample $\omega_{S}$, Eq. (48).